OpenStudy (anonymous):

the equation 5x+2y=6 (a) solve for variable y (b) use this result to find y with x=-4 I think y = 3-5x over 2 (that's a fraction :)) not sure how to do the other part as what I have tried does not make sense thanks in advance

OpenStudy (***[isuru]***):

the first one should be\[\frac{ 6 - 5x }{ 2 }\]not\[\frac{ 3 - 5x }{ 2 }\] ... i think it's a typing mistake... for the second part consider the eqaution\[y = \frac{ 6 - 5x }{ 2 }\] here we don't know the value of x or y.. but the question gives us a value for x which is (-4) and asks as the value of y... all we have to do is find y when x is ( -4) in above equation \[y = \frac{ 6 - 5x }{ 2 }\]\[y = \frac{ 6 - [5\times(-4)] }{ ? }\]\[y = \frac{ 6 - (-20) }{ 2 }\]\[y= \frac{ 6+ 20 }{ 2 }\]\[y= \frac{ 26 }{ 2 }\]\[y= 13\] hope this will help ya!!

OpenStudy (anonymous):

lets hope so and thank you, I will study for a while and hopefully come up with the correct answer :)

OpenStudy (***[isuru]***):

u r welcome!!!